Two charged bodies of charge + 3.0uc and + 7uc and brought in contact for a while and separated.
What are the charges on the bodies now?
1
Expert's answer
2020-09-30T12:26:13-0400

The law of conservation of a charge says

"q_1+q_2=q_1'+q_2'=2q'"

Hence

"q'=\frac{q_1+q_2}{2}=\rm\frac{3.0\:\mu C+7.0\:\mu C}{2}=5.0\:\mu C"